# Maximum Implementable Improvements over Deferred Acceptance

### v2

@@ Abstract
The Deferred Acceptance (DA) mechanism produces stable but Pareto-inefficient school assignments. While mechanisms like Efficiency-Adjusted DA address this inefficiency by asking students to waive priorities at unattainable schools, they can leave significant gains unrealized—sometimes improving only two students when all but one could benefit. Surprisingly, such mechanisms can be doubly dominated, creating fewer improvements while generating more justified envy.
We introduce the Maximum Implementable Improvement over the DA mechanism (MIIDA), which maximizes the number of students made better off given priority-waiving consents. MIIDA weakly Pareto dominates DA, is never doubly dominated, and incentivizes students to consent.
school choice,
maximum improvements,
consent,
C78,
@@
Knipe: tknipe01@qub.ac.uk, Ortega: j.ortega@qub.ac.uk. We acknowledge helpful discussions with Pablo Arribillaga, Péter Biró, Battal Doğan, Martin Meier, and Gabriel Ziegler. T. Knipe acknowledges financial support for doctoral studies from Northern Ireland’s Department for the Economy.
@@ 1 Introduction
The student-proposing Deferred Acceptance (DA) algorithm has become a pillar of market design, combining elegant theoretical properties with widespread adoption in school assignment systems worldwide. While DA guarantees stable matchings, it can produce allocations that are Pareto-inefficient for students in theory and in practice. This inefficiency has motivated the development of mechanisms that weakly Pareto-dominate DA by implementing mutually beneficial trades among students, which only violate the priorities of students who willingly consent to such violations. The most prominent solutions—Efficiency-Adjusted DA (EADA, Kesten 2010) and Top Trading Cycles applied to DA assignments (DA+TTC)—fix DA’s inefficiency without harming any student while, at the same time, preserving weak but meaningful stability and incentive properties. In particular, EADA has shown a remarkable performance in both theory and the lab, and has been considered for practical implementation in Flanders (discussed in Section 2).
In this paper, we analyze how many students can improve their DA assignment through the trading cycles implemented by EADA and DA+TTC, while tracking the number of blocking pairs these exchanges introduce.
@@
We find that both mechanisms can perform poorly when judged solely by the number of students they improve. In problems with $n$ students, we can construct examples where $n-1$ students could benefit. However, EADA and DA+TTC forgo this large cycle involving almost all students and perform a single two-way exchange instead. Therefore, the improvement ratio of both mechanisms, which tells us how many efficiency swaps they miss, is $\frac{n-1}{2}$ (the largest possible ratio obtained by an efficient mechanism that improves upon DA) and grows linearly with the size of the school choice problem (Theorem 1).
Surprisingly, we find that existing mechanisms can be doubly dominated, simultaneously improving fewer students while generating more blocking pairs in specific problems than alternative allocations (Proposition 1). This counterintuitive result reveals an unexploited margin on the efficiency-stability frontier: some Pareto improvements strictly dominate EADA and DA+TTC on both dimensions, yet mechanisms that achieve this have not been explored in the literature. These larger improvements can sometimes require fewer consents overall, and specifically fewer non-beneficiary consents—agreements from students who do not directly benefit from the proposed trades and must therefore act altruistically—compared to the smaller improvements implemented by existing mechanisms (Proposition 2).
Furthermore, EADA and DA+TTC prove incomparable regarding double dominance—in different problem instances, each can produce matchings that doubly dominate the other. This mutual non-dominance suggests that the choice between EADA and DA+TTC involves case-specific tradeoffs rather than universal superiority.
The potentially small number of improvements carried out by existing mechanisms motivates us to propose a new mechanism that always improves the placement of as many students as possible. This algorithm computes DA and, based on the corresponding envy digraph, chooses the collection of disjoint trading cycles that maximizes the number of students involved given the consent structure available. We call this mechanism Maximum Implementable Improvement over Deferred Acceptance (MIIDA). MIIDA weakly Pareto dominates DA and is Pareto-efficient when all students consent. It can generate less or more justified envy, depending on the problem, but it is never doubly dominated because it improves as many students as possible. With regard to incentives, it gives every student a weak incentive to consent to waive their priorities. Finally, it can be computed in polynomial time (Theorem 2).
While MIIDA provides valuable theoretical insights about mechanism performance, it has important drawbacks too: it can be manipulated and can generate more blocking pairs than EADA and DA+TTC. Moreover, the set of blocking pairs that it generates may be a strict superset of the blocking pairs generated by other setwise justified envy minimal mechanisms such as EADA. Such negative properties are unavoidable: any efficient mechanism that weakly Pareto dominates DA and is setwise justified envy minimal has an improvement ratio $\frac{n-1}{2}$, leaving large efficiency gains unrealized (Theorem 3).
Our findings show that consent structures already embedded in existing mechanisms contain unexploited potential for achieving greater welfare improvements, as larger improvements can sometimes be achieved while requiring fewer consents overall, and particularly fewer altruistic consents from non-beneficiary students. Moreover, these larger improvements can occasionally generate fewer blocking pairs than smaller ones, creating a counterintuitive bidimensional improvement on the efficiency-stability trade-off. Overall, MIIDA serves as a useful benchmark that allows us to quantify precisely how much improvement potential is foregone by existing mechanisms and reveals opportunities for further mechanism design in this space.
To summarize, our findings contribute to the literature in three ways. First, we provide tight bounds on the improvement ratios of existing mechanisms, showing that both EADA and DA+TTC can leave significant efficiency gains unrealized in worst-case scenarios, and show that these efficiency gains may be compatible with parallel stability gains. Second, we introduce MIIDA as a new mechanism in the relatively under-explored class of efficient mechanisms that Pareto-dominate DA and document its properties and weaknesses. Third, we establish an impossibility result: any mechanism that attains a non negligible improvement size cannot be justified envy minimal.
@@
The remainder of this paper is organised as follows. Section 2 discusses the literature. Section 3 introduces the model. Section 4 presents the limitations of existing mechanisms. Section 5 introduces the Maximum Implementable Improvement over Deferred Acceptance mechanism. Section 6 concludes.
@@ 2 Related Literature
The Pareto-inefficiency of the student-proposing Deferred Acceptance algorithm has been recognized since the early studies on matching markets and school choice (Roth, 1982; Abdulkadiroğlu and Sönmez, 2003). Kesten (2010) provided the first theoretical bounds on DA’s inefficiency, establishing that the welfare losses can be substantial.
DA’s inefficiency has been consistently documented in practice. Studies including Abdulkadiroglu et al. (2005), Abdulkadiroğlu et al. (2009), Che and Tercieux (2019), Abdulkadiroğlu et al. (2020), and Ortega and Klein (2023) consistently find that DA leaves significant welfare gains on the table across different school choice systems. On the other hand, theoretical literature has identified restrictive conditions on preference and priority structures under which DA achieves efficiency (e.g. Ergin, 2002; Cantillon et al., 2024).
@@
The most prominent solution to DA’s inefficiency is the Efficiency-Adjusted Deferred Acceptance mechanism, proposed in a seminal paper by Kesten (2010). EADA identifies Pareto-improving trading cycles among students that only violate the priorities of students who benefit from consenting to waive their priorities, ensuring such students have no incentive to withhold consent. While Kesten notes that EADA’s improvements may not be the largest possible and argues that selecting larger cycles could be problematic due to the greater number of consents required, our contribution is to quantify exactly how many improvements EADA may be missing and to show that such larger improvements need not always require more consent nor generate more justified envy.
EADA has been extensively studied both theoretically and experimentally, with the literature consistently documenting its strong performance across multiple stability and incentives criteria (Bando, 2014; Tang and Yu, 2014; Dur et al., 2019; Troyan et al., 2020; Troyan and Morrill, 2020; Ehlers and Morrill, 2020; Tang and Zhang, 2021; Reny, 2022; Chen and Möller, 2023; Cerrone et al., 2024; Cheng et al., 2024; Shirakawa, 2025).
Form this literature, the papers closer to ours are Doğan and Ehlers (2021a) and Kwon and Shorrer (2020). These papers show that EADA (among other mechanisms) is justified envy minimal among efficient mechanisms in a set inclusion sense using blocking pairs or triplets, respectively. [^1] Abdulkadiroğlu et al. (2020) and Doğan and Ehlers (2021b) show that the Top Trading Cycles algorithm is also (setwise) justified envy minimal. While this property has made EADA attractive for practical implementation, we prove that any mechanism satisfying setwise justified envy minimality must have an improvement ratio of at least $\frac{n-1}{2}$, leaving large welfare gains unrealized in specific problems. Our MIIDA mechanism deliberately relaxes this constraint to maximize the number of students who benefit, demonstrating that the choice between minimal justified envy and maximal improvement size represents a stark trade-off.
Another natural approach to improving DA is to apply the Top Trading Cycles mechanism to DA’s allocation. While this mechanism is know in the literature, it has received surprisingly limited formal analysis (with some exceptions, see Alcalde and Romero-Medina 2017; Kesten 2010; Troyan et al. 2020).
The broader literature has explored various other approaches to improving DA. Kesten and Kurino (2019) examine conditions under which DA can be improved while maintaining strategy-proofness, while Doğan and Yenmez (2020) study consistency properties of DA improvements. Dur and Morrill (2020) examine the equilibrium performance of mechanisms that dominate DA. Ortega et al. (2025) demonstrate that certain limitations of DA—including rank-inefficiency, inequality and segregation—cannot be addressed by any Pareto-improvement.
@@
Ortega et al. (2024) show that DA is Pareto-inefficient with high probability in large i.i.d. random markets, and that the envy digraph typically contains a giant strongly connected component, implying that most students can potentially benefit from disjoint trading cycles. Furthermore, all efficient mechanisms that dominate DA tend to improve nearly every student. This finding provides context for our analysis. While their paper shows that improvement mechanisms converge in their scope in large i.i.d. random markets, our paper demonstrates that this convergence does not hold in specific instances. We show that significant gaps can emerge between mechanisms in terms of both the number of students improved and the stability costs incurred.
@@ 3 Model
Following Abdulkadiroğlu and Sönmez (2003), a school choice problem $P$ consists of a set of $n$ students $I$ and a set of schools $S$. Each student $i\in I$ has a strict preference relation $\succ_{i}$ over the schools. Each school $s\in S$ has a quota of available seats $q_{s}$ and a strict priority relation $\triangleright_{s}$ over the students. The school $s_{\emptyset}$ with $q_{s_{\emptyset}}=\infty$ denotes being unassigned. For a given school choice problem $P$, a matching $\mu$ is a mapping from $I$ to $S$ such that no school is matched to more students than its quota. We denote by $\mu_{i}$ the school to which student $i$ is assigned.
@@
Student $i$ desires school $s$ in matching $\mu$ if $\textup{rk}_{i}(s)<\textup{rk}_{i}(\mu_{i})$ and he envies student $j$ at matching $\mu$ if $\textup{rk}_{i}(\mu_{j})<\textup{rk}_{i}(\mu_{i})$. We say that student $j$ violates student $i$’s priority at school $s$ in matching $\mu$ if $i$ desires $s$, $\mu_{j}=s$, and $\textup{rk}_{s}(i)<\textup{rk}_{s}(j)$. If so, we call ($i,j,s$) and ($i,s$) a blocking triplet and a blocking pair at matching $\mu$, respectively. A matching $\mu$ is non-wasteful if every school $s$ that is desired by some student in $\mu$ satisfies $|\mu^{-1}_{s}|=q_{s}$. A matching $\mu$ is stable if it is non-wasteful and no student’s priority at any school is violated in $\mu$.
Mechanisms.
A mechanism is a function that maps every school choice problem to a matching. We particularly focus on three mechanisms.
  1. 1.
    Deferred Acceptance (DA). The student-proposing DA mechanism (Gale and Shapley, 1962) works as follows. Students apply to their most preferred school. Schools tentatively accept applicants up to their capacity, rejecting the rest. Rejected students apply to their next choice in subsequent rounds. The algorithm terminates when no new rejections occur, yielding the student-optimal stable matching $\mathtt{DA}(P)$.
  2. 2.
    Efficiency-Adjusted DA (EADA). To execute EADA, first we define an interrupter, which is a student-school pair $(i,s)$ such that:
@@
    To compute EADA, we follow the following procedure:
    1. (a)
      Run DA, find the last interrupter ($i,s$), in which student $i$ consents to waive his priority at school $s$.
    2. (b)
      Eliminate $s$ from the list of preferences of student $i$ and re-run DA.
    3. (c)
      Find the last consenting interrupter. If any, go to step 2. If there are no interrupters, the last DA allocation is implemented.
    When all interrupter students consent to waive their priorities, EADA is Pareto-efficient. In that case, EADA can alternatively be computed by running DA, finding schools that did not reject anyone, matching them permanently to their DA match and removing such students and schools, and re-executing DA on the smaller problem until all schools are rejected (Tang and Yu, 2014).
  3. 3.
    Top Trading Cycles on top of DA (DA+TTC) This mechanism first computes DA, then applies Gale’s Top Trading Cycles algorithm where each student initially "owns" their DA assignment (Shapley and Scarf, 1974). Students point to owners of their most preferred schools, and all students in cycles trade. The process repeats until no cycles remain.
@@
To analyse improvements over DA, we use the envy digraph proposed by Ortega et al. (2024), which illustrates who envies whom after the initial DA allocation (envy need not be justified).
Definition 1.
The envy digraph $G^{\mathtt{DA}(P)}$ is a directed graph where:
  • Nodes represent students in $I$.
  • A directed edge $(i,j)$ exists if student $i$ envies student $j$’s assignment under DA, i.e. if $\textup{rk}_{i}(\mathtt{DA}_{j}(P))<\textup{rk}_{i}(\mathtt{DA}_{i}(P))$.
We denote the edge $(i,j)$ as $i\rightarrow j$. A cycle in the digraph is a sequence of nodes $(i_{0}\to i_{1}\to\ldots\to i_{k}\to i_{0})$ such that there is a directed edge between each pair of consecutive nodes and no edge is repeated. A trading cycle is a cycle in which every node appears exactly once, except for $i_{0}$. Unless specifically stated otherwise, every cycle mentioned henceforth is a trading cycle.
As argued in Section 2, DA is rarely Pareto-efficient. Therefore, we focus on school choice problems where $G^{\mathtt{DA}(P)}$ contains at least one cycle. To systematically identify and implement improvements over DA, we use the concept of a feedback set.
Definition 2.
A feedback set $F$ is a collection of pairwise-disjoint cycles in $G^{\mathtt{DA}(P)}$ such that, when all nodes in these cycles are removed (together with their adjacent edges), the resulting subgraph is acyclic. [^2] Feedback node sets are sometimes called feedback vertex sets (Karp, 2010). The concept of a “cycle cover” is related, but requires that every node must be part of a cycle. The problem of choosing the maximum feedback set is sometimes known as cycle packing.
Feedback sets represent sets of disjoint trading cycles that can be simultaneously implemented to improve upon DA. We use $V(F)$ to denote the set of nodes (students) that are included in some cycle in $F$. If $i\in V(F)$, we say that $F$ covers $i$. A feedback set $F^{\star}$ is maximum if, for any other feedback set $F^{\prime}$, we have $|V(F^{\prime})|\leq|V(F^{\star})|$.
Given a feedback set $F$, we define the matching $\mu^{F}$ as the allocation obtained when the trades in each cycle of $F$ are implemented, starting from DA. Formally:
  • For student $i\in V(F)$ belonging to cycle $C$, $\mu^{F}_{i}=\mathtt{DA}_{\text{succ}(i)}(P)$, where $\text{succ}(i)$ denotes the successor of $i$ in cycle $C\in F$.
  • For student $i\notin V(F)$, $\mu^{F}_{i}=\mathtt{DA}_{i}(P)$.
@@
Every Pareto improvement over DA corresponds to a set of trading cycles in $G^{\mathtt{DA}(P)}$. [^3] Similar arguments have been previously made in the literature by Abdulkadiroğlu et al. (2009); Kesten (2010), Erdil (2014) and Ortega et al. (2024).
Having established the required framework, we are now ready to state our results.

### v3

# The Trade-off Between Minimal Instability and Larger Improvements over Deferred Acceptance

@@ Abstract
The celebrated Efficiency-Adjusted Deferred Acceptance mechanism (EADA) improves the efficiency of the DA algorithm via consented priority violations. Notwithstanding its many merits, we show that EADA can improve only two students when an alternative mechanism that Pareto-dominates DA could benefit all but one student. This shortfall in the number of students improved is not exclusive of EADA but extends to all setwise minimally unstable mechanisms, i.e. those that generate a set of blocking pairs that is never a strict superset of that of another mechanism. The incompatibility between number of students improved and minimal instability disappears when blocking pairs are compared cardinally rather than by set inclusion. In some problems, EADA can be doubly dominated: improving fewer students while generating more blocking pairs.
@@
[…]
Knipe: tknipe01@qub.ac.uk, Ortega: j.ortega@qub.ac.uk. We acknowledge helpful discussions with Péter Biró and Battal Doğan. T. Knipe acknowledges financial support for doctoral studies from Northern Ireland’s Department for the Economy.
@@ 1 Introduction
The student-proposing Deferred Acceptance (DA) algorithm has become a cornerstone of market design, combining elegant theoretical properties with widespread adoption. However, while DA generates stable matchings, it can produce allocations that are Pareto-inefficient for students. The large size of this inefficiency in practice has motivated the development of mechanisms that Pareto-dominate DA by implementing mutually beneficial trades among students, violating only the priorities of students who willingly consent to such violations.111[Abdulkadiroğlu et al. (2009)] report that 4,296 students in NYC would be assigned to a better school without harming any other pupil. The most prominent among these is the Efficiency-Adjusted Deferred Acceptance mechanism (EADA, Kesten 2010), which addresses DA’s inefficiency while preserving weak but meaningful stability and incentive properties. A large literature has shown that EADA performs remarkably well both theoretically and in the lab, making it a key mechanism in the market design literature and leading to its consideration for practical implementation (see Section 2).
In this paper, we study how many students EADA actually improves compared to other mechanisms that also Pareto-dominate DA, and whether these mechanisms improve more students while also generating less instability (as any mechanism that Pareto-dominates DA is bound to create some blocking pairs).
@@
Despite EADA’s excellent performance with regard to other criteria, we find that it can perform poorly when judged solely by the number of students it improves. In school choice problems with $n$ students, we construct examples where n1n-1 students could benefit and yet EADA executes only a single two-way exchange instead (even when all students consent to waive their priorities). Therefore, EADA’s improvement ratio, defined as the size of the largest improvement possible divided by the size of the actual improvement implemented in the worst-case instance, is $\frac{n-1}{2}$, which is the largest possible ratio obtained by any Pareto-efficient mechanism that improves upon DA and grows linearly with $n$ (Theorem 1).
Moreover, the concern that only a few students may improve upon their DA placement is not unique to EADA. In fact, any setwise minimally unstable mechanism that Pareto-dominates DA shares the same limitation (Theorem 2). The impossibility of improving more students while minimizing justified envy in a setwise sense is particularly relevant because each blocking pair represents an unfairly treated student who could potentially challenge the allocation legally. Therefore, mechanism designers who insist on setwise minimizing envy while improving upon DA must accept that the distribution of improvements may be rather limited compared to what other mechanisms can achieve. Conversely, policymakers who intend to maximize the coverage of improvement across the student population must recognize that they may create a strictly setwise larger number of students who are treated unfairly and thus can challenge the assignment.
Taken together, Theorems 1 and 2 extend Kesten’s observation that EADA may not improve as many students as possible by establishing exact bounds on this shortfall and proving an impossibility theorem that highlights the underlying tradeoff between improving more students and being setwise minimally unstable, a property satisfied by EADA among other mechanisms (Doğan and Ehlers, 2021a; Tang and Zhang, 2021).
However, if we compare the instability generated simply in terms of the number of blocking pairs created, we obtain room for improvement in both number of students improved and number of blocking pairs created. We show that EADA can be doubly dominated—simultaneously improving fewer students and generating more blocking pairs than alternative mechanisms that also Pareto-dominate DA (Theorem 3). This previously unknown bi-dimensional possibility of improvement opens an avenue of research into the relatively under-explored class of mechanisms that Pareto-dominate the ubiquitously used Deferred Acceptance mechanism.
@@
The remainder of this paper is organised as follows. Section 2 discusses the literature. Section 3 introduces the model. Section 4 presents our results. Section 5 concludes.
@@ 2 Related Literature
The Pareto-inefficiency of the student-proposing Deferred Acceptance algorithm has been recognized since the early studies on matching markets and school choice (Roth, 1982; Abdulkadiroğlu and Sönmez, 2003). It has been consistently documented in practice (Abdulkadiroglu et al., 2005; Abdulkadiroğlu et al., 2009; Che and Tercieux, 2019; Abdulkadiroğlu et al., 2020; Ortega and Klein, 2023).
@@
The most prominent solution to DA’s inefficiency is the Efficiency-Adjusted Deferred Acceptance mechanism, proposed in a seminal paper by Kesten (2010). EADA identifies Pareto-improving trading cycles among students that only violate the priorities of students who benefit from consenting to waive their priorities, ensuring such students have no incentive to withhold consent. EADA has been extensively studied both theoretically and experimentally, with the literature consistently documenting its strong performance across multiple stability and incentives criteria.
From this literature, the papers closest to ours are Doğan and Ehlers (2021a); Tang and Zhang (2021) and Kwon and Shorrer (2020). These three papers show that EADA is setwise minimally unstable when using blocking pairs or triplets for the comparison, respectively. [^2] Outside of the class of mechanisms that Pareto-dominate DA, Abdulkadiroğlu et al. (2020) and Doğan and Ehlers (2021b) show that the Top Trading Cycles algorithm is also setwise minimally unstable. We contribute to this literature by showing that any setwise minimally unstable mechanism is bound to improve only a few students when many could alternatively benefit.
Furthermore, Doğan and Ehlers (2021a) also show that when comparing minimal instability in a cardinal sense, there is no Pareto-efficient mechanism that Pareto-dominates DA that is cardinally minimally unstable. In this direction, we further show that it is possible to generate fewer blocking pairs than EADA while improving more students.
Beyond these papers, the broader literature has consistently praised EADA’s performance across other dimensions (Bando, 2014; Tang and Yu, 2014; Dur et al., 2019; Troyan et al., 2020; Troyan and Morrill, 2020; Ehlers and Morrill, 2020; Tang and Zhang, 2021; Reny, 2022; Chen and Möller, 2023; Cerrone et al., 2024; Cheng et al., 2024; Dogan and Ehlers, 2023; Shirakawa, 2025).
Another natural approach to Pareto-improve DA is to apply the Top Trading Cycles mechanism to DA’s allocation. While this mechanism is known in the literature, it has received surprisingly limited formal analysis (with some exceptions, see Alcalde and Romero-Medina 2017; Kesten 2010; Troyan et al. 2020). All of our EADA results extend to DA+TTC: it can miss significantly larger improvements while generating more blocking pairs.
More generally, a series of papers have examined when it is possible to improve upon DA while maintaining strategy-proofness (Kesten and Kurino, 2019) or consistency (Doğan and Yenmez, 2020) . However, these Pareto improvements have inherent limits: they cannot address DA’s poor rank distributions or student segregation (Ortega et al., 2025).
@@
Ortega et al. (2024) show that DA is Pareto-inefficient with high probability in large i.i.d. random markets and that most students can potentially benefit from disjoint trading cycles. Furthermore, all efficient mechanisms that dominate DA tend to improve nearly every student. These findings provide context for our analysis. While their paper shows that improvement mechanisms converge in their scope in large i.i.d. random markets, our paper demonstrates that this equivalence does not hold in specific instances, showing that the number of students who benefit can vary dramatically between different efficient mechanisms that dominate DA.
@@ 3 Model
A school choice problem $P$ consists of a set of $n$ students $I$ and a set of schools $S$. Each student $i\in I$ has a strict preference relation $\succ_{i}$ over the schools. Each school $s\in S$ has a quota of available seats $q_{s}$ and a strict priority relation $\triangleright_{s}$ over the students. The school $s_{\emptyset}$ with $q_{s_{\emptyset}}=\infty$ denotes being unassigned. For a given school choice problem $P$, a matching $\mu$ is a mapping from $I$ to $S$ such that no school is matched to more students than its quota. We denote by $\mu_{i}$ the school to which student $i$ is assigned in $\mu$.
@@
Student $i$ desires school $s$ in matching $\mu$ if $\textup{rk}_{i}(s)<\textup{rk}_{i}(\mu_{i})$ and he envies student $j$ at matching $\mu$ if $\textup{rk}_{i}(\mu_{j})<\textup{rk}_{i}(\mu_{i})$. We say that student $j$ violates student $i$’s priority at school $s$ in matching $\mu$ if $i$ desires $s$, $\mu_{j}=s$, and $\textup{rk}_{s}(i)<\textup{rk}_{s}(j)$. In this case, we say that student $i$’s envy towards $j$ is justified, and we call ($i,j,s$) and ($i,s$) a blocking triplet and a blocking pair in matching $\mu$, respectively. We use $B(\mu)$ to denote the set of blocking pairs in matching $\mu$. A matching $\mu$ is non-wasteful if every school $s$ that is desired by some student in $\mu$ satisfies $|\{i\in I:\mu_{i}=s\}|=q_{s}$. A matching $\mu$ is stable if it is non-wasteful and no student’s priority at any school is violated in $\mu$.
Mechanisms with and without Consent. A mechanism is a function that maps every school choice problem to a matching. We focus on two well-known mechanisms. The first of these is the student-proposing Deferred Acceptance mechanism (DA, Gale and Shapley 1962), which works as follows. Students apply to their most preferred school. Schools tentatively accept applicants up to their capacity, rejecting the rest. Rejected students apply to their next choice in subsequent rounds. The algorithm terminates when no new rejections occur, yielding the student-optimal stable matching $\mathtt{DA}(P)$.
The second mechanism of interest is Top Trading Cycles on top of DA (DA+TTC). This mechanism first computes DA, then applies Gale’s Top Trading Cycles algorithm where each student initially “owns” their DA assignment (Shapley and Scarf, 1974). Students point to owners of their most preferred schools, and all students in cycles trade. The process repeats until no cycles remain.
Standard mechanisms simply return a matching for any school choice problem. A consent-based mechanism $M$ instead maps every school choice problem and a subset of students $W\subseteq I$ to a matching. The interpretation of the subset of students $W$ is that these students agree to have their priorities violated at some schools as long as consenting to waive their priorities does not harm them, and thus we refer to $W$ as the consent structure. Formally, the property that waiving their priorities never harms consenting students means that for any problem $P$ and any consent structure $W\subseteq I\setminus\{i\}$:
$\textup{rk}_{i}[M(P,W\cup\{i\})]\leq\textup{rk}_{i}[M(P,W)]$
Note that for any fixed consent structure $W$, a consent-based mechanism $M(\cdot,W)$ reduces to a standard mechanism that maps problems to matchings. Efficiency-Adjusted Deferred Acceptance (EADA, Kesten 2010) is an example of a consent-based mechanism. To describe EADA, we first define an interrupter, which is a student-school pair $(i,s)$ such that:
@@
We say that an interrupter $(i,s)$ consents if $i\in W$. Having defined the notion of an interrupter, we now describe the celebrated (and slightly more complex) EADA consent-based mechanism in detail below.
Algorithm 1 Efficiency-Adjusted Deferred Acceptance (EADA)
1:Input: Problem $P$ and consent structure $W\subseteq I$
2:Compute initial matching $\mu^{0}=\mathtt{DA}(P)$
3:while consenting interrupters exist do
4:  Find the last round $t$ where a consenting interrupter $(i,s)$ was rejected
5:  Identify all consenting interrupters $(i,s)$ in round $t$
6:  if no consenting interrupters found then
7:   break
8:  end if
9:  for each consenting interrupter $(i,s)$ in round $t$ do
10:   Remove school $s$ from student $i$’s preferences
11:  end for
12:  Rerun DA with modified preferences
13:end while
14:return final matching
We denote by $\mathtt{EADA}(P,W)$ the matching generated by Efficiency-Adjusted Deferred Acceptance algorithm in school choice problem $P$ with consent structure $W$. We refer to the mechanism $\mathtt{EADA}(\cdot,I)$ as EADA with full consent.
EADA with full consent can be more straightforwardly computed by identifying every school who did not reject any student in DA’s execution (these schools are called under-demanded). After finding DA’s allocation, those schools are permanently matched with their assigned students, and such under-demanded schools and corresponding students are removed from DA. Then DA is executed in the simplified problem, and this procedure is repeated until all schools have been removed. This procedure was proposed by Tang and Yu (2014) and can be extended to partial consent structures, but this will not be necessary for our purposes.
Setwise and Cardinally Minimally Unstable Mechanisms.
Since it is well-known that every mechanism that Pareto-dominate DA generates blocking pairs in general, we need a way to compare their relative performance in terms of instability (e.g. blocking pairs). We use the notion of setwise minimally unstable mechanisms, a popular way of comparing the blocking pairs generated by assignment mechanisms (Tang and Zhang, 2021; Doğan and Ehlers, 2021a; Abdulkadiroğlu et al., 2020; Kwon and Shorrer, 2020; Bonkoungou and Nesterov, 2025).
Definition 1.
A Pareto-efficient mechanism $M$ is setwise minimally unstable if there does not exist any problem $P$ and any other Pareto-efficient mechanism $N$ such that $B(N(P))\subsetneq B(M(P))$.
EADA with full consent satisfies this property (Doğan and Ehlers, 2021a; Tang and Zhang, 2021), as do many other mechanisms, since the setwise criterion creates only a partial order across mechanisms. Therefore, mechanisms that fail to be setwise minimally unstable are quite disappointing from a legal and fairness perspective, as they generate an unnecessarily large number of students who have their priorities violated and who may legally challenge their allocation. However, as we will demonstrate, this desirable stability property comes at a significant cost in terms of the number of students who can be improved.
A different way to compare the instability among mechanisms that dominate DA is to count the number of blocking pairs they generate, rather than comparing them by set inclusion. This leads to a notion of cardinally more stable mechanisms.
Definition 2.
A Pareto-efficient mechanism $M$ is cardinally more stable than a Pareto-efficient mechanism $N$ if, for every problem $P$, $|B(M(P))|\leq|B(N(P))|$, with strict inequality for some $P$.
We also define two other useful properties of mechanisms used throughout the paper. First, we say that a mechanism $M$ Pareto-dominates DA if: (i) for any problem $P$, the matching $M(P)$ weakly Pareto-dominates the matching $\mathtt{DA}(P)$, and (ii) there exists some problem $P$ where the matching $M(P)$ strictly Pareto-dominates the matching $\mathtt{DA}(P)$. Second, we say that a mechanism $M$ is Pareto-efficient if, for any problem $P$, the matching $M(P)$ is Pareto-efficient.
@@
To analyse improvements over DA, we adopt a graph-theoretic framework. Formally, the envy digraph $G^{\mathtt{DA}(P)}$ is a directed graph where nodes represent students and a directed edge $(i,j)$ exists if student $i$ envies student $j$’s assignment under DA.We denote the edge $(i,j)$ as $i\rightarrow j$.
A cycle $C$ in the envy digraph is a sequence of nodes $C=(i_{0}\to i_{1}\to\ldots\to i_{k}\to i_{0})$ such that there is a directed edge between each pair of consecutive nodes and no edge is repeated. A trading cycle is a cycle in which every node appears exactly once, except for $i_{0}$. Unless specifically stated otherwise, every cycle mentioned henceforth is a trading cycle.
We focus on school choice problems where $G^{\mathtt{DA}(P)}$ contains at least one cycle. A feedback set $F$ is a collection of pairwise-disjoint cycles in $G^{\mathtt{DA}(P)}$ such that, when all nodes in these cycles are removed (together with their adjacent edges), the resulting subgraph is acyclic. [^3] Feedback sets, sometimes called feedback vertex sets, have been extensively studied in graph theory (Karp, 2010). Feedback sets represent sets of disjoint trading cycles that can be simultaneously implemented to improve upon DA.
We use $V(F)$ to denote the set of nodes (students) that are included in some cycle in $F$. If $i\in V(F)$, we say that $F$ covers $i$. Given a feedback set $F$, we define the matching $\mu^{F}$ as the allocation obtained when the trades in each cycle of $F$ are implemented, starting from DA.
@@
Every Pareto improvement over DA corresponds to a set of pairwise-disjoint trading cycles in $G^{\mathtt{DA}(P)}$.
Having established our framework, we are now ready to state our results.